OpenStudy (anonymous):

I need help with this equation please help me work through it! Mario has invested a certain amount of money in stocks and bonds. The annual return on these investments is $730. If the stocks produce $180 more in annual returns than the bonds, how much money does Mario receive per year from each type of investment?

OpenStudy (anonymous):

x+180+x=730, where x is the return from the bonds.

OpenStudy (anonymous):

if you call the amount invested in stocks \(x\) then the amount invested in bonds is \(x-180\) and you know the total is \(730\) so you can solve \[x+x-180=730\] is one way @Numb3r1 has another, where \(x\) is the amount invested in bonds there are many ways to solve

OpenStudy (anonymous):

okay but @satellite73 what do i do after that?

OpenStudy (anonymous):

rewrite as \[2x-180=730\] and solve for \(x\) in two steps 1) add \(180\) to both sides, then 2) divide both sides by \(2\)

OpenStudy (anonymous):

Solve for x and substitute it into the stocks and bonds' values.

OpenStudy (anonymous):

i got \[x= 455\]

OpenStudy (anonymous):

now what should i do?

OpenStudy (anonymous):

@satellite73 can you help me? idk what to do next..

OpenStudy (anonymous):

You used his formula, so x is the larger value and x-180 is the smaller.

OpenStudy (anonymous):

x is the stocks' return, and x-180 is the return from the bonds.

OpenStudy (anonymous):

so 455-180?

OpenStudy (anonymous):

455 from the stocks, what you said from the bonds.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

when i subtracted 455 from 180 i got 275 is that right? @Numb3r1

OpenStudy (anonymous):

You've got it!

OpenStudy (anonymous):

ohhhh okay thank you! :) @Numb3r1